# LSHBloom: Internet-Scale Text Deduplication

### v3

@@ Abstract.
[…]
@@ 1. Introduction
Modern large language model (LLM) training pipelines require the ingestion of many millions of documents to enable internet-scale language pre-training. These documents are obtained through web scraping, deals with academic institutions, and relationships with publishers. The sheer scale of this data ingestion process presents a substantial data management and processing burden when assembling large-scale text databases for continuous LLM pre-training. LLM performance has also been shown to be sensitive to data quality (alignment of data with downstream tasks) (tan202415pintstechnicalreportpretraining) as well as data redundancy–excessive redundancy can cause performance scaling below the predicted rate of established scaling laws (chen-etal-2025-revisiting). Moreover, per-epoch training costs scale in the number of training data points. As such, it is highly desirable to remove outright uninformative data points from internet-scale text datasets–to alleviate the data management burden, to reduce the costs of pre-training, and to enhance LLM performance.
In this work, we focus on reducing data redundancy in text. Duplicate text data–that is, instances of the same content or content that only subtly differ from one another due to differences in publication format, copyediting, and versioning–are ubiquitous in internet-scale text datasets and are well known to lead to a host of undesirable outcomes, from excessive training times to the introduction of predictive biases and training data memorization (lee2021deduplicating), in trained models. Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets (Chandran2024PrivateBT), illegitimately inflating perceived model performance by 8–11% (Dekoninck2024EvadingDC).
These considerations motivate us to address the accurate and efficient deduplication of large numbers (100s of millions) of documents when preprocessing internet-scale text datasets.
We demonstrate via careful evaluation that existing state-of-the-art approaches to text deduplication suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs. Due to the frequent loss of metadata and subtle variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method (broder1998min). The infeasibility of pairwise comparisons in large datasets motivated the development of MinHashLSH (leskovec2014finding), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents. MinHashLSH is by far the most popular method for text deduplication in LLM data pipelines; however, the associated MinHashLSH index can grow extremely large. For example, applying MinHashLSH to the peS2o dataset (peS2o) of just 39M documents takes more than 35 hours on a single 32-core node of the Polaris supercomputer and requires more than 200 GB of disk space (see subsubsection 5.4.1).
@@
MinHashLSH is the most popular and performant state-of-the-art text deduplication technique available today; LSHBloom can be employed as a drop-in replacement for MinHashLSH–preserving this superior performance while enabling significant gains in scaling capacity as we increase our training dataset sizes. We know from historical data that LLM training dataset sizes will almost quadruple each year (epoch2024datasetsizetrend). LSHBloom presents an opportunity to create robust data processing infrastructure for LLM pre-training that is effective, efficient, and scalable well into the future. In summary, the contributions of our work are the following:
@@
    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data
@@
    We present the first study on the effects of tuning the hyperparameters for each of these algorithms, and present the best parameter settings for each
@@
    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime
@@ 2. Background
We first provide background information on the deduplication problem and describe the current state-of-the-art for large-scale text deduplication pipelines.
2.1. Defining the Deduplication Problem
Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus. These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document). Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”). In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization. High computational performance is critical due to the extreme size of modern LLM datasets. For example, peS2o (peS2o) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature (gusenbauer2019google), DOLMA (soldaini_dolma_2024) contains roughly 5B documents, and RedPajama (together2023redpajama) has just over 100B documents—scales at which naive pairwise comparisons become infeasible. Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
Here, we focus on methods for document-level deduplication. Suppose an operator $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1. We define our task as follows: Given a corpus of $N$ documents, $D_{N}=$ $d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, identify every document $d$ such that $\exists d^{\prime}\in D_{N}$ such that $d^{\prime}\neq d$ and $S(d,d^{\prime})\geq T$.
@@ 2.1. Defining the Deduplication Problem
Just how to define duplication is a subtle question. For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above. However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will induce small differences even when applied to the same underlying document. Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
For these reasons, many contemporary deduplication techniques take a ”fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs. One way to do this is to alter one’s choice of similarity metric to encapsulate a broader notion of similarity. This is common in so-called semantic deduplication techniques such as SemDeDup (abbas2023semdedup), which employ embeddings from pre-trained language models to identify semantically similar text data. Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents. It is more efficient to instead establish a similarity metric based on the raw text content of each document.
2.2. Jaccard Similarity and MinHash Approximation
Popular techniques such as those used to build the DOLMA (soldaini_dolma_2024) and DataComp-LM (li2024datacomp) datasets measure n-gram overlap between documents to identify duplicates. Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
@@ 2.2. Jaccard Similarity and MinHash Approximation
In this paradigm, a text document is viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
@@
However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming. The MinHashing procedure (broder1998min) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (broder1998min; leskovec2014finding). Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity. This estimate grows more accurate as we increase our sample size.
2.3. MinHashLSH for Scalable MinHash
@@
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of documents for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
@@
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH (leskovec2014finding) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This procedure enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
@@
One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents. Additionally, inserting and querying these tree-structured indices constitutes a significant portion (¿85%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1). Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
@@
Figure 1. Breakdown of wall clock time on 10% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
@@ 2.4. Bloom Filter
The Bloom filter (bloom1970space) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
A Bloom filter consists of a bit array of size $m$, with all bits initially set to 0. $k$ independent hash functions are defined, each of which hashes an element to an integer in the range 1 to $m$. To insert an element into the Bloom filter, it is passed through the $k$ hash functions, and the resulting $k$ positions in the bit array are set to 1. We can then check if an element is in the set by hashing using the same $k$ hash functions. If all $k$ of the corresponding positions in the bit array are 1, the filter returns a positive. If any of those $k$ bits is 0, the element is not contained in the set.
Bloom filters have been used extensively in distributed computing for a variety of purposes (broder2004network; chervenak2008globus; tarkoma2011theory).

### v4

@@ Abstract.
Modern internet-scale text databases require robust entity deduplication to maintain data integrity and reduce processing overhead. While Locality-Sensitive Hashing (LSH) is the established paradigm for near-duplicate detection, traditional LSHIndex structures based on trees or hashmaps suffer from large disk footprints and pointer-chasing latencies. We propose LSHBloom, a novel architectural extension to MinHashLSH that replaces traditional indexing structures with an array of independent probabilistic Bloom filters. By mapping LSH bands directly to contiguous bit vectors, optimizing for spatial locality, and co-designing our software with modern supercomputer hardware, LSHBloom achieves $12\times$ higher throughput than standard MinHashLSH on the 39M-document peS2o dataset. Our design reduces disk footprint by $18\times$ while maintaining near-identical recall and precision. We provide a formal analysis of the analytically boundable error rates and demonstrate, through extrapolation, that LSHBloom enables billion-scale record deduplication on commodity hardware.
@@ 1. Introduction
Modern text databases must handle numerous texts at the internet-scale, particularly with the advent of data-hungry systems like large language models (LLMs) that require access to many millions of documents. Redundant text records are a major source of inefficiency in these systems—recent studies have demonstrated that large web-scraped corpora such as CommonCrawl (Nagel, 2023) (roughly 300B webpages) are comprised of between 14% to 52% near-duplicate records (Fröbe et al., 2021). Many modern applications that rely on text databases benefit strongly from procedures that mitigate data redundancy including 1) internet-scale LLM pre-training where data redundancy has marked negative effects on language modeling ability (Tan and Wang, 2024; Lee et al., 2021) and significantly inflates training costs (Chen et al., 2025), 2) retrieval-augmented generation systems that utilize vector databases where storing $O(d)$ fewer documents results in $O(log(D))$ faster query times (Malkov and Yashunin, 2018) and mitigates the substantial GPU cost involved in embedding text for insertion (Gao et al., 2023), 3) full-text search systems where storing $O(d)$ fewer documents results in a linear decrease in storage costs and in query complexity (Zobel and Moffat, 2006; Ayyagiri et al., 2024), 4) database tasks such as topic modeling or clustering where $O(d)$ fewer documents translates into an $O(d)$ speedup (Blei et al., 2003; Grootendorst, 2022), and 5) lookup tasks, such as those in search engines, that benefit from operations that can group text data by near-duplicates to group or distill query results (e.g., near-duplicate detection, fuzzy grouping) (Charikar, 2002; Andoni et al., 2018; Broder, 1997).
Traditional relational entity resolution techniques based on exact-matching or common field values (e.g., blocking) (Christen, 2012) fall short for text due to parsing noise (aberrations in raw text content introduced when parsing from PDF, XML, HTML, etc. into raw text) and the frequent absence or unreliability of metadata when obtaining documents from the web. Consequently, recent work has focused on the use of approximate similarity metrics such as n-gram overlap (Soldaini and others, 2024; Li et al., 2024) and the Jaccard similarity (Broder, 1997; Broder et al., 1998).
The computational infeasibility of performing pairwise comparisons between documents in large text datasets motivated the development of MinHashLSH (Leskovec et al., 2014), which applies a locality-sensitive hashing (LSH) scheme to avoid comparing dissimilar documents. MinHashLSH is by far the most popular method for text deduplication in LLM data ingestion workflows (Brown et al., 2020; Rae et al., 2021; Olmo et al., 2025; Touvron et al., 2023; Dubey and others, 2024; Computer, 2023; Gao et al., 2020; Mireshghallah et al., 2023; Black et al., 2022; Biderman et al., 2023; Penedo et al., 2025, 2023; Li et al., 2023; Lozhkov et al., 2024); however, the associated MinHashLSH index can become slow and grow extremely large when applied at internet scale, presenting bottlenecks in throughput as well as in memory and disk resources. For example, applying MinHashLSH to the tiny peS2o dataset (Soldaini and Lo, 2023) containing 39M academic documents takes more than 35 core hours on the Polaris supercomputer and requires more than 200 GB of disk space (see Section 5.4.1). Overcoming these bottlenecks is paramount: language models are routinely retrained with newly acquired data (Dai et al., 2024; Parmar et al., 2024) and vector databases are continually updated with new content (Taipalus, 2024; Ma et al., 2023). Continuously updating these collections requires efficient data ingestion, cleaning, and deduplication workflows. Moreover, when constructing text datasets, upstream changes in data pipelines such as the mode of data ingestion, cleaning, or preprocessing may alter the representation of existing documents and therefore, necessitate rerunning deduplication workflows on large corpora to mitigate data redundancy. Beyond a certain size, the runtime and disk requirements imposed by existing state-of-the-art deduplication techniques, such as MinHashLSH, render these processes infeasible to run.
@@
LSHBloom is designed as a drop-in replacement for the industry-standard MinHashLSH (Brown et al., 2020; Olmo et al., 2025; Dubey and others, 2024; Rae et al., 2021)—preserving deduplication fidelity while providing significant headroom in scaling capacity as the cardinality of text datasets increases. Historical analysis of frontier models indicates that LLM training dataset sizes almost quadruple each year (Rahman and Owen, 2024). Beyond the increasing availability of large-scale general-purpose web corpora (Gao et al., 2020; Penedo et al., 2023, 2025) (including CommonCrawl (Nagel, 2023), which adds 3-5B newly scraped web pages each month), coordinated national initiatives such as the Genesis Mission (Executive Office of the President, 2025) will integrate decades of scientific research into new, large-scale datasets for LLM training. This trend towards larger text datasets is further exacerbated by the recent strategic shift towards using high-volume synthetic data generation to continue scaling LLM training (Gunasekar et al., 2023; Javaheripi, 2023) despite limitations in the amount of available human-authored data. The ease with which one can generate vast quantities of synthetic text contributes to an ever-increasing data management burden. These observations motivate the development of LSHBloom, a robust data deduplication technique for large text datasets that is effective, efficient, and scalable well into the future. In summary, our contributions are:
@@
    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized by LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data.
@@
    We present an early study on the effects of tuning the hyperparameters for each of these algorithms, and identify the best parameter settings for each.
@@
    We carefully co-design the implementation with the modern supercomputer hardware to maximize throughput.
  4. (4)
    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime.
@@ 2. Background
We first introduce the deduplication problem and describe the state-of-the-art for large-scale text deduplication pipelines.
2.1. Defining the Deduplication Problem for Text
In the context of database systems, text deduplication is a specialized form of Online Duplicate Elimination. Unlike batch-oriented entity resolution, which materializes all pairwise similarities within a static table (Christen, 2012), text deduplication in modern ingestion pipelines functions as a Streaming Approximate Membership Query (SAMQ). The objective is to determine, in real-time, if an incoming record represents a redundant entry relative to the existing corpus state.
Formally, we define this task as a set-similarity membership problem. Suppose a similarity function $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1. We define our task as follows: Given a stream of $N$ documents, $D_{N}=$ $d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, for each document, $d_{i}$, evaluate the binary decision function $\mathcal{F}(d_{i})$ which returns 1 if $\exists d_{j}\in D_{seen}:S(d_{i},d_{j})\geq T$ where $D_{seen}=\{d_{j}\mid j<i\}$ and 0 otherwise. This formulation prioritizes ingestion throughput and state compression as storing the full history, $D_{seen}$, is intractable at scale.
@@ 2.1. Defining the Deduplication Problem for Text
The choice of similarity function, $S$, is crucial for maximizing join fidelity. For example, one could simply stipulate that two documents must be byte-for-byte identical. However, this definition may exclude many document pairs that are not exact matches, but have high overlap in their content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. Also, when web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will produce small differences in raw text content even when applied to the same underlying document. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective. As a result, properly deduplicating text data requires adopting approximate similarity measures that are robust to minor aberrations.
2.2. Approximate Similarity Metrics for Text
One way to define an approximate similarity measure between documents is to measure the proportion of n-grams in that document that were previously seen in the data stream. Popular techniques such as those used to build the DOLMA (Soldaini and others, 2024) dataset and DataComp-LM (Li et al., 2024) baseline dataset employ this procedure to identify duplicates. However, these techniques can be sensitive to the relative frequency of n-grams within a document or dataset.
@@ 2.2. Approximate Similarity Metrics for Text
Alternatively, some techniques, such as MinHashLSH, take a set-theoretic perspective on document similarity. In this set-theoretic paradigm, a text document is viewed as a set of n-grams, and the similarity between two such sets, $A$ and $B$, is defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
@@
However, computing the Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For extremely high-dimensional objects such as documents, computing the Jaccard similarity explicitly can be extremely expensive. The MinHashing procedure (Broder et al., 1998) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures—lossy, fixed-length representations of the underlying high-dimensional objects. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (Broder et al., 1998; Leskovec et al., 2014). Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity. This estimate grows more accurate as we increase our sample size (the number of random permutations).
2.3. MinHashLSH for Text Deduplication
@@
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of MinHash signatures for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
@@
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH (Leskovec et al., 2014) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This procedure enables us to produce a series of candidate pairs of documents that we can either take as duplicate pairs (our approach in this work) or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces errors due to the use of locality-sensitive hashing.
@@
One can implement a prefix-tree or a hashmap to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents. Moreover, navigating these data structures may require pointer-chasing and, therefore, non-contiguous memory accesses that cause cache misses. Inserting into and querying against these indices constitutes a significant portion (¿85%) of algorithm runtime, acting as the primary bottleneck for real-time stream throughput at extreme scales (see Figure 1). Here, we remedy this space and throughput issue by employing a novel combination of MinHashLSH and Bloom filters.
@@
Figure 1. Breakdown of wall clock time on 10% of peS2o for conventional MinHashLSH and our LSHBloom method
@@ 2.4. Bloom Filter
The Bloom filter (Bloom, 1970) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
A Bloom filter consists of a contiguous bit array of size $m$, with all bits initially set to 0. $k$ independent hash functions are defined, each of which hashes an element to an integer in the range 1 to $m$. To insert an element into the Bloom filter, it is passed through the $k$ hash functions, and the resulting $k$ positions in the bit array are set to 1. We can then check if an element is in the set by hashing using the same $k$ hash functions. If all $k$ of the corresponding positions in the bit array are 1, the filter returns a positive. If any of those $k$ bits is 0, the element is not contained in the set.
Bloom filters have been used extensively in distributed computing for a variety of purposes (Broder and Mitzenmacher, 2004; Chervenak et al., 2008; Tarkoma et al., 2011).